# J1: Incentivizing Thinking in LLM-as-a-Judge via Reinforcement Learning

### v1

@@ 4 Results
4.1 Main Results on Benchmarks with Pairwise-J1
@@
Comparison of J1 with state-of-the-art methods on PPE. Table 1 compares our Pairwise-J1 models with different baselines on the PPE benchmark. J1-Llama-70B, our larger model, obtains an overall accuracy of $69.6$, outperforming all previous methods, including those trained on much more data (see column 2). Compared to related competing approaches that are generative reward models, J1-Llama-70B outperforms both EvalPlanner (Saha et al., 2025) by $1.7\%$ ($67.9\rightarrow 69.6$), and DeepSeek-GRM-27B (Liu et al., 2025b) by $7.1\%$ ($62.2\rightarrow 69.3$). First, this shows the effectiveness of J1’s training methodology and use of online RL (GRPO), compared to EvalPlanner, which is trained on the same data but with two iterations of DPO. Second, this also highlights the effectiveness of J1’s high-quality synthetic preference pairs, compared to the data used to train DeepSeek-GRM-27B. The latter is first fine-tuned on 1270K judge data, followed by stages of Reinforcement Learning on 237K samples and further scaling at test time with a meta reward model across 32 generations. J1 shows particularly strong results on the PPE Correctness subset that spans popular reasoning benchmarks, thereby highlighting its utility as a reward model for Best-of-N alignment. At a smaller scale, J1-Llama-8B is competitive with Armo-8B (scalar RM) and outperforms EvalPlanner-8B and a larger Skywork-Reward-Gemma-2-27B by significant margins ($54.1\rightarrow 59.8$ and $55.6\rightarrow 59.8$, respectively).
Test-time Scaling of J1. We further scale J1 up at test time with techniques such as self-consistency (Wang et al., 2023). As shown in Table 1 (see “w/ SC@32” rows), we obtain further improvements of up to $1.5\%$ at both scales of J1 by sampling 32 CoTs (with a temperature of $1.0$) and using self-consistency to obtain the final verdict. In subsection 4.2, we conduct a more detailed analysis of test-time scaling of J1.
Comparison of J1 with Thinking-LLMs. Given the superiority of the earlier Thinking-LLM method EvalPlanner over other baselines, Table 2 compares J1 to EvalPlanner on all other benchmarks at both scales (8B/70B). Additionally, we compare J1 to general Thinking-LLMs (e.g., DeepSeek-R1-Distill-Llama, DeepSeek-R1, and OpenAI-o1-mini). DeepSeek-R1-Distilled-Llama is SFT-ed with 800K long CoTs from DeepSeek-R1 (a much larger 671B MoE model), starting from the same base models as J1. Thus, in a head-to-head comparison where the base models are the same, J1 outperforms DeepSeek-R1-Distilled-Llama across all benchmarks by large margins. In particular, on the harder JudgeBench benchmark that requires evaluating GPT-4o responses, J1-Llama-70B obtains an absolute 14% improvement ($46.0\rightarrow 60.0$) over R1-Distilled-Llama-70B. J1 also closes the gap to the much larger R1 model, even outperforming it on RewardBench by 2.7%.
To further analyze this observation, Table 8 in Appendix C shows a detailed breakdown of the results for the individual categories in RewardBench. Although R1 expectedly excels in reasoning, J1 outperforms it in evaluating non-verifiable prompts like those in Chat-Hard and Safety. Overall, our results suggest that J1 is a generalist judge that can be used for evaluating responses to both verifiable and non-verifiable prompt tasks, and in different stages of the LLM development process.
Models Type PPE Correctness JudgeBench
(a, b) (b, a) Consistent Verdict- (a, b) (b, a) Consistent Verdict-
Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Flip/Ties $\downarrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Flip/Ties $\downarrow$
Llama-3.1-8B-Instruct Pairwise 54.7 54.1 30.2 44.1 67.4 42.3 32.3 37.4
EvalPlanner-Llama-8B Pairwise 53.0 52.4 37.3 25.7 38.2 62.8 30.3 30.5
J1-Llama-8B (random single-order data) Pairwise 58.3 57.6 38.3 36.7 48.3 59.4 36.6 32.9
J1-Llama-8B (both-order data) Pairwise 59.2 58.4 39.1 36.8 63.1 51.4 42.0 27.7
J1-Llama-8B (verdict consistency reward) Pairwise 58.4 58.2 43.9 28.7 52.3 64.6 45.4 26.0
J1-Llama-70B (both-order data) Pairwise 72.9 72.3 61.2 21.9 68.9 72.6 60.0 20.6
Llama-3.1-8B-Instruct Pointwise 42.9 25.1 46.3 26.9
J1-Llama-8B Pointwise 53.8 9.4 57.7 8.0
J1-Llama-70B Pointwise 65.0 13.7 60.0 16.6
@@
Table 3: Results on PPE Correctness and JudgeBench, comparing various position-bias mitigation strategies (for Pairwise-J1) and Pointwise-J1. Consistent Accuracy measures the proportion of examples where the pairwise judge gives the correct verdict in both response orders. Verdict Flip shows the percentage of cases where a pairwise judge’s verdict flips when the order of responses is swapped, regardless of correctness. Ties indicate the proportion of examples where a pointwise judge assigns equal scores to both responses. All the generations are with greedy decoding.
4.2 Ablations and Analyses of Pairwise-J1 and Pointwise-J1
@@
In Section 2.5, we introduced several alternative training recipes for Thinking-LLM-as-a-Judge models, including the Pointwise-J1 variant. In the following experiments, we systematically analyze the impact of these recipes, focusing on three key research questions: (i) How does Pointwise-J1 compare to Pairwise-J1 in mitigating position bias? (ii) How can we design effective rewards for reinforcement learning in Thinking-LLM-as-a-Judge models? (iii) What is the impact of different thinking prompts on model behavior? Unless otherwise specified, we use Pairwise-J1 to refer to the model trained under our primary Pairwise with Verdict setting. Other variants are discussed explicitly where relevant.
Figure 4: Test-time scaling of Pairwise-J1 and Pointwise-J1 on the PPE Correctness benchmark. We show greedy decoding results in dotted lines for comparison. As we sample more $N$, (i) position-consistent accuracy increases and (ii) tie rate decreases, for both pairwise and pointwise models, at both 8B and 70B scales.
Pairwise-J1 vs Pointwise-J1 and their Effects on Positional Consistency. There are two main ways of mitigating position bias in judgments – either by improving a pairwise judge itself (e.g., by training on both orders of data, adding consistency-based rewards, etc., like we do in J1) or by training a Pointwise-J1 model that, by design, is position-consistent. Pointwise judges, while consistent, are score-based and thus could suffer from ties if both responses receive the exact same score.
To understand this long-standing issue of judgment bias, in Table 3, we report (i) individual accuracy for both orders of responses, (ii) position-consistent accuracy, and (iii) verdict-flips/ties. We use “verdict flip” and “tie” to refer to the same metric which is the fraction of samples where either the verdict changes based on the response order (for pairwise judges) or both responses are scored equally (for pointwise judges). A consistently correct judge is thus expected to show higher position-consistent accuracy and lower verdict-flips/ties. We observe that when judgment quality is measured by a stricter position-consistency accuracy, Pointwise-J1 outperforms Pairwise-J1. The same trend holds between Pointwise and Pairwise-Llama-Instruct as well. While our bias mitigation techniques on Pairwise-J1 help, our best Pairwise-J1 model still suffers from a significant fraction of inconsistencies (up to 20%) whereas for Pointwise-J1, the fraction of ties is as low as 9%, leading to improvement in position-consistent accuracy. Scaling up our Pointwise-J1 model to 70B improves position-consistent accuracy further.
@@
Figure 5: Distribution of Absolute Scores and $\Delta Score\leavevmode\nobreak\ (Chosen-Rejected)$ generated by the 8B Pairwise-J1 (w/ Scores) and Pointwise-J1 models on the PPE Correctness benchmark. Pairwise-J1 exhibits sparser score distribution and larger score differences between Chosen and Rejected (ground-truth) responses.
Pairwise-J1 8B Variants PPE Preference PPE Correctness RewardBench JudgeBench RM-Bench FollowBenchEval
@@
Positive Reward for Correct Verdict 60.3 59.2 85.7 42.0 73.4 48.3
+ Negative Reward for Incorrect Verdict 60.0 59.1 85.4 44.9 70.8 42.0
+ Format Reward 60.2 58.3 85.6 40.3 71.8 49.3
@@
Thinking (default - Figure 8) 60.3 59.2 85.7 42.0 73.4 48.3
Plan & Execution (Figure 7) 59.1 58.9 85.8 44.3 71.8 50.2
Effect of Test-time Scaling on Position-Consistent Accuracy and Ties. In Table 1, we showed results with one kind of test-time scaling of J1 – self-consistency over multiple verdicts. For a Pointwise-J1 model, we can also scale it up at test-time by averaging scores across multiple generations to get a more accurate estimate of these scores. In Figure 4, we plot position-consistent accuracy and ties as a function of number of generations $N$ (refer to Table 7 in Appendix C for accuracy numbers). We observe improved position-consistent accuracy and reduction in ties for both Pairwise- and Pointwise-J1 at both 8B and 70B scales. To the best of our knowledge, this is one of the first comprehensive analyses of Pointwise versus Pairwise judges, when both are trained using the same data.
Effect of Reward Schemes and Seed Prompts for Training J1. In Table 4, we first study the effect of different rewards for Pairwise-J1 models. We obtain best results when only assigning positive rewards to correct verdicts – adding additional format rewards or negative rewards for incorrect verdicts marginally degrades performance. Next, we also analyze the effect of two different seed prompts that are used to elicit “thinking” in J1 models. Our default J1 Thinking prompt is motivated by DeepSeek-R1. Additionally, similar to EvalPlanner (Saha et al., 2025), we experiment with a prompt that instructs the model to first plan for the evaluation recipe, then execute the evaluation according to that recipe and the response(s), before generating the final verdict (Figure 8 in Appendix A). We find that J1 is robust to such choices, performing comparably with both prompts. In fact, with a simpler Thinking prompt, the model tends to generate richer reasoning traces, including evaluation criteria, reference answers, re-evaluations, and detailed comparisons (see Figure 2).
Score Distribution of Pointwise-J1 and Pairwise-J1. In Figure 5, we show the distribution of (i) absolute scores and (ii) score difference between the chosen and rejected (ground-truth) responses, as generated by Pairwise-J1 w/ Scores and Pointwise-J1 models. Pairwise-J1 exhibits a sparser distribution of scores (clustered around a few scores) and a larger gap compared to Pointwise-J1. This difference arises from their training objectives: Pairwise-J1 directly contrasts both responses within the same input, allowing it to more clearly differentiate between them, which is reflected in the score assignments. In contrast, Pointwise-J1 is trained with distant supervision that attends to only one response at a time, making it harder to learn fine-grained comparative judgments.
Reward and Thought Length Analysis of J1. Figure 6 illustrates the training and validation reward, as well as average generation length throughout different stages of J1 training. As training progresses, the thought lengths of most pairwise judges converge at around 500 tokens, while the pointwise judge tends to generate shorter outputs, typically between 300 and 400 tokens due to the absence of comparison-style tokens. Training rewards for the pairwise variants exhibit a similar steady increase before converging. In Table 5 of Appendix C, we provide a comparison of different Pairwise-J1 variants, all of which show comparable performance.

### v2

@@ 4 Results
4.1 Main Results on Benchmarks with Pairwise-J1 and MultiTask-J1
Results on PPE Correctness. We first evaluate J1 on PPE Correctness because it directly tests RMs and LLM-Judges for their ability to improve popular reasoning benchmarks. This benchmark also offers a broader potential for improvement compared to others, such as RewardBench. Table 1 shows the results. J1-Qwen3-32B-MultiTask, our best J1 model, obtains state-of-the-art performance with an overall accuracy of $76.8$, outperforming all previous methods by significantly large margins ($p<0.0001$), including those trained on much more data (see column 2). Compared to related competing approaches that are generative reward models, J1-Qwen3-32B-MultiTask outperforms both EvalPlanner (Saha et al., 2025) by $6.8\%$ ($70.2\rightarrow 76.6$), and DeepSeek-GRM-27B (Liu et al., 2025b) by $17\%$ ($59.8\rightarrow 76.8$). J1 also improves upon the base Qwen3-32B model, a strong thinking-LLM, by a large $10.3\%$ ($66.5\rightarrow 76.8$).
@@ 4.1 Main Results on Benchmarks with Pairwise-J1 and MultiTask-J1
Furthermore, all J1 models at three different scales outperform their base counterparts, highlighting the generalizability of our recipe. In particular, this validates the effectiveness of J1’s training methodology and use of online RL, compared to EvalPlanner, which is trained on the same data but with two iterations of DPO. Second, this also shows the effectiveness of J1’s high-quality synthetic preference pairs, compared to the data used to train DeepSeek-GRM-27B. The latter is first fine-tuned on 1270K judge data, followed by stages of Reinforcement Learninng on 237K samples and further scaling at test time with a meta reward model across 32 generations. At a smaller scale, J1-Llama-8B is competitive with Armo-8B (scalar RM) and outperforms EvalPlanner-8B and a larger Skywork-Reward-Gemma-2-27B by significant margins ($52.8\rightarrow 59.2$ and $54.7\rightarrow 59.2$, respectively).
Table 2: Results on five reward modeling benchmarks, where PPE includes both correctness and preference subsets of data. We compare J1 at different scales to EvalPlanner and general Thinking-LLMs (distilled-R1, o1-mini, o3, and R1). We report the default metric for each benchmark, where $\dagger$: JudgeBench and FollowBenchEval use positional consistent accuracy. Reported gains are relative to the corresponding base models.
Models Overall PPE RewardBench RM-Bench JudgeBench FollowBenchEval
Llama-3.1-8B-Instruct   0048.3 55.6   00069.5   00054.0   000032.3   0000030.2
DeepSeek-R1-Distilled-Llama-8B   0054.7 58.9   00073.7   00069.3   000030.5   0000040.9
EvalPlanner-Llama-8B   0056.2 54.3   00083.0   00068.1   000030.2   0000045.3
J1-Llama-8B   0061.9+13.6 59.8 +4.2   00085.7 +16.2   00073.4 +19.4   000042.0 +9.7   0000048.3 +18.1
Llama-3.3-70B-Instruct   0064.3 65.8   00085.4   00069.5   000048.6   0000052.2
DeepSeek-R1-Distilled-Llama-70B   0067.4 68.6   00086.9   00080.8   000046.0   0000054.6
EvalPlanner-Llama-70B   0073.2 67.9   00093.8   00082.1   000056.6   0000065.4
J1-Llama-70B   0075.0+10.7 69.6 +3.8   00093.3 +7.9   00082.7 +13.2   000060.0 +11.4   0000069.3 +17.1
Qwen3-32B   0077.3 66.5   00090.9   00088.1   000070.8   0000070.0
J1-Qwen-32B-MultiTask   0080.8+3.5 71.8 +5.1   00093.6 +2.7   00090.3 +2.2   000071.4 +0.6   0000077.1 +7.1
OpenAI-o1-mini   0072.7 68.5   00087.1   00080.8   000064.2   0000062.9
OpenAI-o3   0077.4 72.1   00086.4   00086.1   000075.7   0000066.8
DeepSeek-R1-671B   0078.4 72.3   00090.6   00088.6   000068.9   0000071.7
@@
Table 3: Results on PPE Correctness comparing judgment bias of pairwise and pointwise J1. Consistent Accuracy: proportion of examples judged correctly in both response orders. Verdict Flip: proportion of cases where the pairwise verdict changes when order is swapped. Ties: proportion of examples where the pointwise judge assigns equal scores.
Models Type (a, b) Acc $\uparrow$ (b, a) Acc $\uparrow$ Consistent Acc $\uparrow$ Verdict Flip/Ties $\downarrow$
J1-Llama-70B Pairwise 72.9 72.3 61.2 21.9
J1-Llama-70B Pointwise 65.0 13.7
J1-Qwen-32B Pairwise 74.6 74.2 65.2 14.5
J1-Qwen-32B Pointwise 69.3 13.0
J1-Qwen-32B-MultiTask Pairwise 76.8 76.2 67.0 17.0
J1-Qwen-32B-MultiTask Pointwise 70.6 10.5
Results on RewardBench. Table 6 in Appendix D shows a comparison of J1 with leading generative reward models on RewardBench. J1-Qwen-32B-MultiTask obtains an overall score $93.6$, outperforming all previous methods. Importantly, J1 is equally performant on all four categories of RewardBench. This suggests that J1 is a generalist judge that can be used for evaluating responses to both verifiable and non-verifiable prompt tasks, and in different stages of the LLM development process.
Comparison of J1 with Thinking-LLMs. Next, in Table 2, we compare J1 to several SOTA Thinking-LLMs on all benchmarks at three different scales (8B/32B/70B). These include DeepSeek-R1-Distill-Llama, DeepSeek-R1, Qwen3-32B, OpenAI-o1-mini, and OpenAI-o3. DeepSeek-R1-Distilled-Llama is SFT-ed with 800K long CoTs from DeepSeek-R1 (a much larger 671B MoE model), starting from the same base models as J1. Thus, in a head-to-head comparison where the base models are the same, J1 Llama models outperform DeepSeek-R1-Distilled-Llama across all benchmarks by large margins. Impressively, J1-Qwen-32B-MultiTask also outperforms state-of-the-art thinking models, R1 and o3, on three out of five benchmarks, while only training a 32B model with synthetic data. Through these comprehensive evaluations, J1 thus reinforces the utility of online RL and synthetic data for training state-of-the-art Thinking-LLM-as-a-Judge models.
4.2 Ablations and Analyses of J1 in Pointwise vs. Pairwise Setups
@@
In this section, we systematically analyze the impact of different J1 formulations introduced in §2.3, focusing on following research questions: (i) How does pointwise setup compare to pairwise in mitigating position bias? (ii) What benefit does multitasking bring compared to separate Pointwise-J1 and Pairwise-J1 training? (iii) How can we design effective rewards for reinforcement learning in Thinking-LLM-as-a-Judge models? (iv) What is the impact of different thinking prompts on model behavior?
@@
Figure 3: Distribution of Absolute Scores and $\Delta Score~(Chosen-Rejected)$ on PPE Correctness from J1-Qwen-32B-Multitask when used in pairwise and pointwise settings. Pairwise exhibits sparser score distribution and larger score differences between Chosen and Rejected (ground-truth) responses. Note that all samples with a positive $\Delta$ in the range [0, 1) are correct predictions and count toward the ‘0’ bar in the plot.
Figure 4: Test-time scaling of J1 on PPE Correctness. We show greedy decoding results in dotted lines for comparison. As we sample more NN, (i) position-consistent accuracy increases and (ii) tie rate decreases, for both pairwise and pointwise models, at both 8B and 70B scales.
Position Consistency. There are two main ways of mitigating position bias in judgments – either by improving a pairwise judge itself (e.g., by training on both orders of data, adding consistency-based rewards, etc., like we do in J1) or by training a Pointwise-J1 model that, by design, is position-consistent. While on one hand, pairwise judges can be position-inconsistent, pointwise judges on the other hand are consistent but lack the context of reference candidates to ground their evaluations on and hence are more prone to ties.
To understand this long-standing issue of judgment bias, in Table 3, we report (i) individual accuracy for both orders of responses, (ii) position-consistent accuracy, and (iii) verdict-flips/ties. We use “verdict flip” and “tie” to refer to the same metric which is the fraction of samples where either the verdict changes based on the response order (for pairwise judges) or both responses are scored equally (for pointwise judges). A consistently correct judge is thus expected to show higher position-consistent accuracy and lower verdict-flips/ties.
We observe that when judgment quality is measured by a stricter position-consistent accuracy, Pointwise-J1 outperforms Pairwise-J1. However, if we consider a random ordering of response pairs, Pairwise-J1 performs better. Moreover, the multitask formulation leverages the strengths of both pointwise and pairwise judges, outperforming separately trained judges in both random-order and position-consistency accuracy. In Table 8 and Appendix D, we also study the effect of training with both orders of data and verdict consistency rewards on Pairwise-J1 models.
Pairwise-J1 8B Variants Overall PPE RewardBench JudgeBench RM-Bench FollowBenchEval
@@
Positive Reward for Correct Verdict 61.8 59.8 85.7 42.0 73.4 48.3
+ Negative Reward for Incorrect Verdict 60.4 59.6 85.4 44.9 70.8 42.0
+ Format Reward 61.0 59.3 85.6 40.3 71.8 49.3
@@
Thinking (default - Figure 10) 61.8 59.8 85.7 42.0 73.4 48.3
Plan & Execution (Figure 9) 62.1 59.0 85.8 44.3 71.8 50.2
Score Distribution of Pointwise vs Pairwise Evaluation. Recall that J1-Qwen-32B-MultiTask has the advantage of being evaluated in both pointwise and pairwise settings. To understand the difference in scores assigned by the model in both these judgment settings, Figure 3 shows the distribution of (i) absolute scores and (ii) score differences between the chosen and rejected (ground-truth) responses. Both settings assign high scores more frequently (e.g., 8–10 are the tallest bars), but pairwise exhibits a sparser distribution and a larger gap between chosen and rejected responses, compared to pointwise. This difference stems from their training objectives: pairwise directly contrasts both responses within the same input, enabling clearer differentiation reflected in the scores. In contrast, pointwise is trained with distant supervision on one response at a time, making fine-grained comparative judgments harder. Therefore, combining both approaches can leverage their complementary strengths for improved judgment quality.
Effect of Test-time Scaling. We explore test-time scaling of J1 by either conducting self-consistency over multiple verdicts or averaging over multiple pointwise scores. We show that these can lead to further improvements (see Table 12 and 13 in Appendix D). Here, we specifically analyze whether J1 models that generate scores can achieve more accurate judgments and fewer ties by averaging scores across multiple generations at test time. In Figure 4, we plot position-consistent accuracy and ties as a function of number of generations NN. We observe improved position-consistent accuracy and reduction in ties for both Pairwise- and Pointwise-J1 at both 8B and 70B scales. To the best of our knowledge, this is one of the first comprehensive analyses of Pointwise versus Pairwise judges, when both are trained using the same data.
@@
Effect of Reward Schemes and Seed Prompts for Training J1. In Table 4, we first study the effect of different rewards for Pairwise-J1 models. We obtain best results when only assigning positive rewards to correct verdicts – adding additional format rewards or negative rewards for incorrect verdicts marginally degrades performance. Next, we also analyze the effect of two different seed prompts that are used to elicit “thinking” in J1 models. Our default J1 Thinking prompt is motivated by DeepSeek-R1. Additionally, similar to EvalPlanner (Saha et al., 2025), we experiment with a prompt that instructs the model to first plan for the evaluation recipe, then execute the evaluation according to that recipe and the response(s), before generating the final verdict (Figure 10 in Appendix B). We find that J1 is robust to such choices, performing comparably with both prompts. In fact, with a simpler Thinking prompt, the model tends to generate richer reasoning traces, including evaluation criteria, reference answers, re-evaluations, and detailed comparisons (see Figure 1).
Reward and Thought Length Analysis of J1. Figure 5 illustrates the training and validation reward, as well as average generation length throughout different stages of J1 training. As training progresses, the thought lengths of most pairwise judges converge at around 500 tokens, while the pointwise judge tends to generate shorter outputs, typically between 300 and 400 tokens due to the absence of comparison-style tokens. Training rewards for the pairwise variants exhibit a similar steady increase before converging. In Table 7 of Appendix D, we provide a comparison of different Pairwise-J1 variants, all of which show comparable performance.
Example 1 of J1’s Thinking Trace for a Verifiable Math Question <think>
Okay, let’s see. The user wants to know which assistant’s answer is better. The problem is about finding the sum of the sums of all two-element subsets of the set {1,2,3,4,5,6}.
First, I need to check if both answers are correct. The correct answer should be 105. Let me verify that. The number of two-element subsets in a 6-element set is C(6,2) = 15, which matches the problem statement. Each element appears in (6-1) = 5 subsets. So each number from 1 to 6 is included 5 times in the total sum. The sum of the numbers 1 through 6 is 21. So total sum would be 5*21 = 105. That’s the correct answer.
Now, looking at Assistant A’s answer. They list all 15 subsets, compute each sum correctly, add them up step by step, and get 105. The final answer is boxed correctly. The steps are clear and methodical.
Assistant B’s answer also lists the subsets and computes each sum. However, when adding them up, they make a mistake. The sums they list are correct, but when they add them, they group them incorrectly. For example, they say 3 + 4 + 5 + 6 + 7 + 5 + 6 + 7 + 8 + 7 + 8 + 9 + 9 + 10 + 11. Let me add those numbers: 3+4=7, +5=12, +6=18, +7=25, +5=30, +6=36, +7=43, +8=51, +7=58, +8=66, +9=75, +9=84, +10=94, +11=105. Wait, that’s 105. But Assistant B says they simplified to 47. That’s a clear error. They must have miscalculated when adding. The final answer they give is 47, which is wrong.
So Assistant A’s answer is correct, while Assistant B’s is incorrect. The user’s problem requires the correct answer, so Assistant A’s response is better. Additionally, Assistant B’s explanation of the addition step is flawed. They mention grouping the sums into 13 + 10 + 11 + 13 = 47, which doesn’t make sense. The correct total is 105.
Therefore, Assistant A’s answer is accurate and follows the instructions properly, while Assistant B’s answer contains a critical error leading to an incorrect result. The user’s instructions emphasized providing a clear, step-by-step solution, which Assistant A did correctly. Assistant B’s answer, despite being detailed, has a mistake in the final addition step, leading to an incorrect answer.
</think>
Figure 6: Thinking trace of J1 for a verifiable prompt. J1 first self-generates a reference answer to the math problem (highlighted in brown). Subsequently, it checks the correctness of Assistant A’s answer and judges it as correct (highlighted in green). Then it concludes that Assistant B’s answer is incorrect and also provides feedback by pointing to a calculation mistake (highlighted in blue).
Qualitative Analysis of J1’s Thinking Traces. Figure 6 shows J1’s thinking trace for a math prompt. We see that the model starts thinking by first self-generating a reference answer. Thereafter, it judges that Assistant A’s response is correct, and for Assistant B’s response, it analyzes each of the steps and identifies a calculation mistake. It then summarizes the evaluation by providing critical feedback of how and where to improve Assistant B’s response. Figure 7 and 8 in Appendix A show two more examples of J1’s thinking for a verifiable math prompt and a non-verifiable writing prompt.